OpenStudy (anonymous):

Use the iterative rule to find the 7th term in the sequence. an = 30 – 4n a7 = _________

OpenStudy (studygurl14):

Just plug in 7 into the equation and solve. One question, is it 4n or \(4^n\) ?

OpenStudy (anonymous):

I can't paste it but your second 4n.

OpenStudy (studygurl14):

Okay. So just do this: \(\large a_n=30-4^2\rightarrow a_7=30-4^7\) Do you have a calculator that can do exponents?

OpenStudy (anonymous):

Not with me.

OpenStudy (studygurl14):

Use this: http://www.calculatorsoup.com/calculators/algebra/exponent.php

OpenStudy (studygurl14):

remember: \(\large -4^7=-(4)^7\)

OpenStudy (studygurl14):

@Shawn.Wilkerson u there?

OpenStudy (anonymous):

Yes, I'm still working it out. I was AFK for a few seconds.

OpenStudy (studygurl14):

okay. What's "AFK"?

OpenStudy (perl):

away from keyboard

OpenStudy (studygurl14):

ah. thanks @perl